The firm Dread Pirate Roberts, Inc. produces pirate costumes as well as pirate movies. The product pirate costumes have $61,000 in revenue, and pirate movies have $30,000 revenue.

Costs for pirate costumes are $71,000, and costs for pirate movies are $10,000.

If the firm drops pirate costumes, then revenue for pirate movies will decrease by 15%. The firm cannot avoid $1,000 of the cost of producing pirate costumes.

More than one downstream retailer may go out of business if the Dread Pirate Roberts, Inc. does not carry pirate costumes because these products are an essential part of these distributors' value proposition.

a.

Relevant cost analysis suggests it is $5,500 MORE profitable to the drop pirate costumes product than to keep it. Strategic cost analysis, though, suggests that the firm should keep it.

b.

Relevant cost analysis suggests it is $4,500 LESS profitable to the drop pirate costumes product than to keep it. Strategic cost analysis, though, suggests that the firm should drop it.

c.

Relevant cost analysis suggests it is $4,500 MORE profitable to the drop pirate costumes product than to keep it. Strategic cost analysis, though, suggests that the firm should keep it.

d.

Relevant cost analysis suggests it is $5,500 LESS profitable to the drop pirate costumes product than to keep it. Strategic cost analysis, though, suggests that the firm should drop it.
Answer #1

If Dread pirates Roberts Inc. drops pirate costumes it will save 71,000 - 61,000 = 10,000 as opportunity gain,

also it will lose 15% of 30,000 = 4,500

and it cannot avoid 1000,

Net benefit will be =10,000-4500-1000 = 4,500, hence,

c. Relevant cost analysis suggests it is $4,500 MORE profitable to the drop pirate costumes product than to keep it. Strategic cost analysis, though, suggests that the firm should keep it.